1.      Use Gaussian elimination to solve the system of linear equations


300x1 112x2 109x3 = 521

252x1 156x2 330x3 =738

108x1 -123x2 121x3 =106

2.     Solve the following system linear equations by Gauss Jordan Method


x +y +z = 5

2x +3y +5z = 8

4x + 5z = 2


1
Expert's answer
2021-11-29T16:08:32-0500

1.

"\begin{pmatrix}
 300 & 112&109&|&521 \\
 252 & 156&330&|&738 \\
108 & -123&121&|&106 \\
\end{pmatrix}"


divide the 1 row by 300:


"\begin{pmatrix}
 1 & 28/75&109/300&|&521/300 \\
 252 & 156&330&|&738 \\
108 & -123&121&|&106 \\
\end{pmatrix}"


multiply 1 row by 252 and subtract it from 2 row; multiply 1 row by 108 and subtract it from 3 row:


"\begin{pmatrix}
 1 & 28/75&109/300&|&521/300 \\
 0 & 61.92&238.44&|&300.36 \\
0 & -163.32&81.76&|&-81.56 \\
\end{pmatrix}"


divide the 2 row by 61.92:


"\begin{pmatrix}
 1 & 28/75&109/300&|&521/300 \\
 0 & 1&1987/516&|&2503/516 \\
0 & -163.32&81.76&|&-81.56 \\
\end{pmatrix}"


multiply 2 row by 28/75 and subtract it from 1 row; (multiply 2 row by 163.32 and add it to 3 row:


"\begin{pmatrix}
 1 & 0&-1663/1548&|&-115/1548 \\
 0 & 1&1987/516&|&2503/516 \\
0 &0&122235/172&|&122235/172 \\
\end{pmatrix}"


divide the 3 row by 122235/172:


"\begin{pmatrix}
 1 & 0&-1663/1548&|&-115/1548 \\
 0 & 1&1987/516&|&2503/516 \\
0 &0&1&|&1 \\
\end{pmatrix}"


(multiply 3 row by 1663/1548 and add it to 1 row; multiply 3 row by 1987/516 and subtract it from 2 row:


"\begin{pmatrix}
 1 & 0&0&|&1 \\
 0 & 1&0&|&1 \\
0 &0&1&|&1 \\
\end{pmatrix}"


"x_1=x_2=x_3=1"


2.

"\begin{pmatrix}
 1 & 1&1&|&5 \\
 2 & 3&5&|&8 \\
4 & 0&5&|&2 \\
\end{pmatrix}"


multiply 1 row by 2 and subtract it from 2 row; multiply 1 row by 4 and subtract it from 3 row:


"\begin{pmatrix}
 1 & 1&1&|&5 \\
 0 & 1&3&|&-2 \\
0 & -4&1&|&-18 \\
\end{pmatrix}"


multiply 2 row by 1 and subtract it from 1 row; multiply 2 row by 4 and add it to 3 row:


"\begin{pmatrix}
 1 & 0&-2&|&7 \\
 0 & 1&3&|&-2 \\
0 & 0&13&|&-26 \\
\end{pmatrix}"


divide the 3 row by 13:


"\begin{pmatrix}
 1 & 0&-2&|&7 \\
 0 & 1&3&|&-2 \\
0 & 0&1&|&-2 \\
\end{pmatrix}"


multiply 3 row by 2 and add it to 1 row; multiply 3 row by 3 and subtract it from 2 row:


"\begin{pmatrix}
 1 & 0&0&|&3 \\
 0 & 1&0&|&4 \\
0 & 0&1&|&-2 \\
\end{pmatrix}"


"x=3,y=4,z=-2"